# Dynamic Attention-Guided Context Decoding for Mitigating Context Faithfulness Hallucinations in Large Language Models

### v1

@@ B.2 Training and Evaluation
Appendix C Details of LR Classifier Generalization and Feature Analysis
C.1 Cross-Prompt Test Results
To evaluate the robustness of the classifier under different prompts, we reconstructed the attention ratio feature vectors using Prompts 2, 3, and 4. These prompts differ in structure and phrasing but are consistent in task objectives. The classifier, trained using Prompt 1, was then tested on these alternative prompts.
@@ C.1 Cross-Prompt Test Results
C.2 Analysis of the Importance of Different Features
@@
Appendix D Analyzing Context Utilization Signal Contribution Across Heads
D.1 Signal Strength and Concentration in Heads
@@
D.2 Complementary Contributions of Weaker Heads
@@
D.3 Data-Dependent and Generalizable Heads
To investigate whether the influential heads identified in one dataset maintain their significance across other datasets, we conducted cross-dataset experiments. For datasets $A$, $B$, and $C$, we analyzed the top 10 heads ranked by LR coefficients and observed significant differences in head rankings across datasets. However, when considering the intersection of the top 100 heads across all three datasets, we identified a subset of heads that consistently exhibited high importance.
Using this subset of generalizable heads, we retrained the classifier on dataset $A$ and tested it on datasets $B$ and $C$. The model achieved high performance across all datasets, suggesting the presence of data-independent heads that encode universal contextual utilization signals. Conversely, dataset-specific top heads exhibited superior performance on their respective datasets, indicating a degree of data dependency in some heads’ contributions.
@@ D.3 Data-Dependent and Generalizable Heads
  3. 3.
    Data Dependency and Generalization: Some heads exhibit data-specific importance, while others demonstrate generalizability across datasets, reflecting a balance between dataset-specific and universal contributions.
@@
Appendix E Models and Datasets
E.1 Model Details
  • LLaMA2-7B: meta-llama/Llama-2-7b-hf
  • LLaMA2-7B-Chat: meta-llama/Llama-2-7b-chat-hf
  • LLaMA2-13B: meta-llama/Llama-2-13b-hf
  • LLaMA2-13B-Chat: meta-llama/Llama-2-13b-chat-hf
  • Mistral-7B: mistralai/Mistral-7B-v0.1
  • Mistral-7B-Instruct: mistralai/Mistral-7B-Instruct-v0.1
Appendix F Experimental Details
@@
F.1 Dataset Categories and Statistics
Multi-Hop Reasoning (HotpotQA). HotpotQA Yang et al. (2018) is a benchmark dataset for multi-hop reasoning, requiring models to synthesize information across multiple paragraphs to generate an answer. This dataset emphasizes complex reasoning over distributed evidence, making it a critical benchmark for evaluating context utilization.
Long-Form Retrieval-Based QA (TriviaQA, SearchQA). TriviaQA Joshi et al. (2017) and SearchQA Dunn et al. (2017) require reasoning over longer contexts, with answers scattered across retrieved documents. These datasets test the model’s ability to focus on relevant content in lengthy contexts and generate precise answers.
Single-Paragraph Extraction (SQuAD). SQuAD Rajpurkar et al. (2016), and NewsQA Trischler et al. (2017) are standard extractive QA datasets where the answer is typically located within a single paragraph. These datasets are widely used for evaluating the span-extraction capabilities of QA systems.
Document-Level QA (NQ). NQ Kwiatkowski et al. (2019) is a document-level open-domain question answering dataset driven by real user queries. It requires systems to extract long answers from entire Wikipedia documents and generate specific short answers, evaluating document-level information retrieval and natural language understanding capabilities.
Simulated Conflict Scenarios (NQ-Swap). NQ-Swap Longpre et al. (2021) is an artificially constructed dataset that introduces adversarial entity swaps into NQ to create ambiguous or conflicting contexts. It evaluates the model’s ability to resolve conflicts and faithfully utilize context.
F.2 Dataset Sources and Formats
@@ F.3 Dataset Statistics
F.4 Baseline Method Configurations
We implemented the baseline methods with their recommended hyperparameter settings for fair comparisons:
@@ F.4 Baseline Method Configurations
F.5 Implementation Details
All experiments utilized a unified prompt template (Prompt 1) to ensure consistency across methods. The prompt format is detailed in Appendix G. For decoding, greedy decoding was employed to produce deterministic outputs and facilitate direct comparisons across methods. All models run on NVIDIA A100 GPUs.
@@ F.5 Implementation Details
LLaMA2-7B-Chat Regular 0.2368 0.9846 0.9056 0.8035 0.8515
@@
F.6 Results on Summarization Tasks
@@
F.7 Additional Results of The Impact of Different $\alpha$
@@
F.8 Impact of Different Prompts
@@
Appendix G Prompt Templates
@@
Appendix H Symbol Explanation
The symbols and corresponding meanings defined in this paper, as shown in Table 7.
Symbol Meaning
$H_{\text{norm}}(P)$ Normalized entropy of distribution $P$
$v_{j}$ Attention ratio feature vector for token $j$
$v_{j}^{(K)}$ Top-K attention ratio feature vector for token $j$
$a_{l,h}^{j}$ The raw attention of the $h$-th head in the $l$-th layer for token $j$
$r_{l,h}^{j}$ The attention ratio of the $h$-th head in the $l$-th layer for token $j$
$s_{j}$ The utilization score for token $j$
$c_{k}$ The coefficient of the $k$-th feature in the LR classifier
$w_{k}$ The normalized coefficient of the $k$-th feature in the LR classifier
$U$ The utilization distribution
$u_{i}$ The normalized utilization score for token $j$
$U_{top}$ The subset of $U$ for tokens ranked in the top-$R$ of the generation distribution
Table 7: Annotation table.

### v2

@@ B.1 LR Classifier Training and Evaluation
B.2 Cross Prompt Templates Testing
To evaluate the robustness of the classifier under different prompts, we reconstructed the attention ratio feature vectors using Prompts 2, 3, and 4. These prompts differ in structure and phrasing but are consistent in task objectives. The classifier, trained using Prompt 1, was then tested on these alternative prompts. (Templates shown in Figure 15)
@@ B.2 Cross Prompt Templates Testing
Appendix C Analyzing Context Utilization Signal Contribution Across Heads
C.1 Analysis of the Importance of Different Features
@@
C.2 Signal Strength and Concentration in Heads
@@
C.3 Complementary Contributions of Weaker Heads
@@
Appendix D Experimental Details
D.1 Dataset Details
We conducted experiments on seven open-book question-answering (QA) datasets, representing a variety of QA tasks. These include multi-hop reasoning datasets (HotpotQA Yang et al. (2018)), long-form retrieval-based QA datasets (TriviaQA Joshi et al. (2017) and SearchQA Dunn et al. (2017)), single-paragraph extraction tasks (SQuAD Rajpurkar et al. (2016) and NewsQA Trischler et al. (2017)), and document-level QA datasets (NQ Kwiatkowski et al. (2019)). All datasets are formatted in the unified schema provided by the MrQA repository Fisch et al. (2019). Additionally, we used the artificially constructed NQ-swap dataset Longpre et al. (2021), designed to simulate conflicting or ambiguous scenarios by replacing entities.
D.1.1 Dataset Categories and Statistics
@@
Multi-Hop Reasoning (HotpotQA).
HotpotQA Yang et al. (2018) is a benchmark dataset for multi-hop reasoning, requiring models to synthesize information across multiple paragraphs to generate an answer. This dataset emphasizes complex reasoning over distributed evidence, making it a critical benchmark for evaluating context utilization.
Long-Form Retrieval-Based QA (TriviaQA, SearchQA).
TriviaQA Joshi et al. (2017) and SearchQA Dunn et al. (2017) require reasoning over longer contexts, with answers scattered across retrieved documents. These datasets test the model’s ability to focus on relevant content in lengthy contexts and generate precise answers.
Single-Paragraph Extraction (SQuAD, NewsQA).
SQuAD Rajpurkar et al. (2016), and NewsQA Trischler et al. (2017) are standard extractive QA datasets where the answer is typically located within a single paragraph. These datasets are widely used for evaluating the span-extraction capabilities of QA systems.
Document-Level QA (NQ).
NQ Kwiatkowski et al. (2019) is a document-level open-domain question answering dataset driven by real user queries. It requires systems to extract long answers from entire Wikipedia documents and generate specific short answers, evaluating document-level information retrieval and natural language understanding capabilities.
Simulated Conflict Scenarios (NQ-Swap).
NQ-Swap Longpre et al. (2021) is an artificially constructed dataset that introduces adversarial entity swaps into NQ to create ambiguous or conflicting contexts. It evaluates the model’s ability to resolve conflicts and faithfully utilize context.
D.1.2 Dataset Sources and Formats
@@ D.1.3 Dataset Statistics
D.2 Implementation Details
[…]
D.3 Model Details
  • LLaMA2-7B: meta-llama/Llama-2-7b-hf
  • LLaMA2-7B-Chat: meta-llama/Llama-2-7b-chat-hf
  • LLaMA2-13B: meta-llama/Llama-2-13b-hf
  • LLaMA2-13B-Chat: meta-llama/Llama-2-13b-chat-hf
  • Mistral-7B: mistralai/Mistral-7B-v0.1
  • Mistral-7B-Instruct: mistralai/Mistral-7B-Instruct-v0.1
D.4 Baseline Configurations
We compare the proposed method DAGCD with three decoding strategies: Greedy Decoding, CAD Shi et al. (2024b), and COIECD Yuan et al. (2024). CAD and COIECD are specifically designed to mitigate context faithfulness hallucination. We implemented the baseline methods with their recommended hyperparameter settings for fair comparisons:
@@ D.4 Baseline Configurations
LLaMA2-7B-Chat Greedy 0.2368 0.9846 0.9056 0.8035 0.8515
@@
D.5 Results on Summarization Tasks
@@
Appendix E Detailed Results of "5.3 Ablation Study"
E.1 Additional Results of "Ablation 1: Detector Training Data Sizes"
Table 12 shows the performance variations of all LLMs used in this study when trained with detectors on different amounts of data. As observed, our method maintains consistent performance across various data sizes for all LLMs.
Figure 12: Detector Training Data Size Validation. The variation in inference performance across different models when using detectors trained on varying amounts of data.
E.2 Additional Results of "Ablation 3: Scaling Factor $\alpha$"
@@
E.3 Impact of Different Prompts
@@
Appendix F Prompt Templates